x+y+z=6, x-y+z=2, 2x+3y-z=5
In these system of equations use graphing, elimination and substitution

Answer by Theo(13342)

if you can eliminate one of the variables, then you might be able to solve using two dimensional graphing software for two of the variables.

after that, it's a simple matter to solve for the third.

it's probably best to eliminate the z if you can because most two dimensional graphing software work with x and y only.

your equations are:

x + y + z = 6 (equation 1)
x - y + z = 2 (equation 2)
2x + 3y - z = 5 equation 3)

add equation 1 to equation 3 to get:

3x + 4y = 11 (equation 4)

subtract equation 2 from equation 1 to get:

2y = 4 (equation 5)

equations 4 and 5 can be graphed to find their intersection.

the intersection of those 2 equations on the graph is (1,2).

that means that x = 1 and y = 2

use those values for y and go back to any of the original equations to find that z = 3.

all original equations are true when x = 1 and y = 2 and z = 3, confirming that the solution is correct.

here's the graph of equations 4 and 5.

those equations are:

3x + 4y = 11 (equation 4)
2y = 4 (equation 5)

$$$